OpenStudy (anonymous):

write the expression 27^-2/3 in radical notation. evaluate the radical expression

OpenStudy (tkhunny):

One piece at a time. \[27^{-2/3} = \frac{1}{27^{2/3}} = \frac{1}{\left(\sqrt[3]{27}\right)^{2}} = \frac{1}{3^{2}}\] One more thing to go!